# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

### v3

@@
Abstract
The study of time series is crucial for understanding trends and anomalies over time, enabling predictive insights across various sectors. Spatio-temporal data, on the other hand, is vital for analyzing phenomena in both space and time, providing a dynamic perspective on complex system interactions. Recently, diffusion models have seen widespread application in time series and spatio-temporal data mining. Not only do they enhance the generative and inferential capabilities for sequential and temporal data, but they also extend to other downstream tasks. In this survey, we comprehensively and thoroughly review the use of diffusion models in time series and spatio-temporal data, categorizing them by model category, task type, data modality, and practical application domain. In detail, we categorize diffusion models into unconditioned and conditioned types and discuss time series and spatio-temporal data separately. Unconditioned models, which operate unsupervised, are subdivided into probability-based and score-based models, serving predictive and generative tasks such as forecasting, anomaly detection, classification, and imputation. Conditioned models, on the other hand, utilize extra information to enhance performance and are similarly divided for both predictive and generative tasks. Our survey extensively covers their application in various fields, including healthcare, recommendation, climate, energy, audio, and transportation, providing a foundational understanding of how these models analyze and generate data. Through this structured overview, we aim to provide researchers and practitioners with a comprehensive understanding of diffusion models for time series and spatio-temporal data analysis, aiming to direct future innovations and applications by addressing traditional challenges and exploring innovative solutions within the diffusion model framework.
Index Terms:
@@ Abstract
survey, diffusion model, time series, spatio-temporal data, generative model, temporal data, DDPM, Score SDE.
1 Introduction
@@
Diffusion models represent a family of probabilistic generative models that undergo optimization through a two-step process involving the injection and subsequent removal of noise across a set of training samples. This process comprises a forward phase, referred to as diffusion, and a reverse phase, known as denoising. By training the model to remove the noise added during the diffusion process, the model learns to generate effective data samples during inference that align closely with the distribution of the training data [2, 189].
@@
Figure 1: Trends in the cumulative number of papers related to diffusion models for time series and spatio-temporal data.
Figure 2: An overview of diffusion models for time series and spatio-temporal data analysis. In diffusion process, $x_{k}$ and $x_{k-1}$ denote the results after adding noise at step $k$ and $k-1$, respectively. This process can be represented by the size of the controlling steps $\beta_{k}\in(0,1)$, the identity matrix $\mathbf{I}$, and a Gaussian distribution $\mathcal{N}(x;\mu,\sigma)$ of $x$ with the mean $\mu$ and the covariance $\sigma$. During the denoising process, the model attempts to iteratively learn the data distribution by modelling the distribution ${p}_{\theta}(x_{k-1}|x_{k})$. The functions $\mu_{\theta}(\cdot)$ and variance $\sigma_{\theta}(\cdot)$ are the model learnable parameters.
@@
In recent years, diffusion models have risen to prominence and significantly influenced various domains, including computer vision (CV) [2, 8, 190, 191, 192], natural language processing (NLP) [193, 195, 172, 194], and general multimodal learning [196, 112, 197, 198]. This challenges the long-time supremacy of generative adversarial networks (GANs) [187, 115]. Within these areas, diffusion models have demonstrated remarkable capabilities in applications such as text-to-image [112, 199], instance segmentation [200, 201], 3D shape generation [203, 202], molecule design [204, 205, 292], and audio generation [206, 91]. Remarkably, diffusion models have also gained popularity as a non-autoregressive alternative for tasks conventionally dominated by autoregressive methods [189]. Recently, the introduction of OpenAI Sora [87] marks the advent of diffusion models in modeling the physical world embedded within the spacetime continuum, highlighting their critical importance. In addition, AlphaFold 3 [292] proposed by Google DeepMind uses diffusion models to generate 3D atomic coordinates and predict biomolecular structures like proteins, DNA, and RNA.
@@
Temporal data, which primarily includes time series and spatio-temporal data, encapsulates the dynamics of the vast majority of real-world systems [76]. These forms of temporal data have been extensively studied and are recognized as crucial for numerous applications [77, 207, 208]. However, deriving universal dynamic laws in the physical world from various data modalities remains a significant challenge within the field. Recently, the area of time series and spatio-temporal modeling has experienced a substantial shift from sensory intelligence towards general intelligence [209]. This shift is characterized by the emergence of unified foundation models (FMs) that possess versatile temporal data analytical capabilities [76, 209], challenging the supremacy of domain-specific models. Diffusion models have achieved state-of-the-art results on many modalities, including images, speech, and video [210]. Benefiting from the vast and diverse available data in these fields, diffusion models often serve as generative FMs alongside large language models (LLMs) or other foundation models, facilitating rapid development in these areas [7, 190]. In recent years, there has also been an increasing number of diffusion models crafted for modeling time series and spatio-temporal data (Figure 1 ). Also, we have become aware of an increasing number of attempts to use diffusion models for temporal modeling (see Table I ). Observing the success of diffusion models, an intriguing question arises: what kind of sparks will emerge from the intersection of time series/spatio-temporal data analysis and diffusion models?
@@
Time series and spatio-temporal data analysis fundamentally rely on a profound understanding of their inherent temporal dynamics, wherein primary tasks predominantly focus on the generative capabilities of backbone models, such as forecasting [13, 86, 253], imputation [56, 262, 69] and generation [182, 198]. These analyses center on generating temporal data samples for specific purposes in conditional or unconditional manners. Having witnessed the recent development of time series and spatio-temporal foundation models [76, 254], whether built upon LLMs or trained from scratch, their success can be attributed to the ability to estimate the distribution of training samples where effective data representations can be drawn. In this regard, diffusion models emerge as a powerful generative framework that enables (1) the modeling of complex patterns within temporal data and (2) the support of a wide range of downstream tasks, as depicted in Figure 2 .
@@
To generate valid data samples for specific tasks, time series and spatio-temporal diffusion models usually operate in an unconditional manner without the need for supervision signals. Given the partially-observed nature of real-world applications [255], conditional diffusion models have emerged. They leverage data labels (e.g., instructions, metadata, or exogenous variables) to regulate the generation process, thereby enabling effective cross-modal prompting that leads to more tailored and improved outcomes [182]. We present a roadmap in Figure 3 . By training on large-scale temporal data, diffusion models effectively fill the gap of time series/spatio-temporal data generation and exhibit significant potential in solving the puzzle of next-generation, LLM-empowered temporal data-centric agents [256, 209].
@@
Figure 3: Representative diffusion models for time series and spatio-temporal data in recent years.
@@
Despite the promising prospects and rapid advancement of diffusion models in handling time series and spatio-temporal data, there has been a conspicuous lack of systematic analysis of this model family in the existing literature. This article aims to bridge this gap by providing a forward-looking review that elucidates both the ’why’ and the ’how’ —- detailing the reasons diffusion models are suited for these data modalities and unveiling the mechanisms through which they confer advantages. In this survey, we offer a detailed categorization, engage in thorough reviews, and identify burgeoning trends within this rapidly evolving landscape. Our main contributions are summarized as follows:
@@
2 Background
This paper primarily reviews recent improvements in using diffusion models to solve time series and spatio-temporal data challenges. In this section, we will first define time series and spatio-temporal data, as well as their corresponding tasks in various fields. Then, we will introduce the history of diffusion model and its advantages. Finally, some different kinds of diffusion models and their variants based on theoretical formula derivations and comparisons with other generative models will be presented.
2.1 Overview of Time Series and Spatio-Temporal Data
Temporal data, particularly time series and spatio-temporal data, are important data structures for a wide range of real-world applications [76]. A time series is defined as a sequential arrangement of data points, categorized by their temporal order. These sequences may be univariate, involving a single variable over time, or multivariate, incorporating multiple variables. For instance, daily air quality measurements in a city constitute a univariate time series, while combining daily temperature and humidity readings generates a multivariate series. In our discussions, we employ bold uppercase letters (e.g., $\mathbf{X}$) to represent matrices, bold lowercase (e.g., $\mathbf{x}$) for vectors, calligraphic uppercase (e.g., $\mathcal{X}$) for sets, and standard lowercase (e.g., $x$) for scalars.
We base our formal definitions of time series on those provided in [76, 77]. Specifically, a univariate time series (Figure 4 (a)), denoted as $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T}$, consists of a sequence of $T$ data points arranged chronologically, where each $x_{t}\in\mathbb{R}$ represents the series’ value at time $t$. Conversely, a multivariate time series (Figure 4 (b)), represented by $\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T})\in\mathbb{R}^{T\times D}$, encompasses a sequence of $T$ data points also in temporal sequence but across $D$ feature channels, with $\mathbf{x}_{t}\in\mathbb{R}^{D}(1\leq t\leq T)$ indicating the series’ values at time $t$ across $D$ different channels. For a comprehensive exploration of time series, we direct the reader to [77].
@@ 2.1 Overview of Time Series and Spatio-Temporal Data
Figure 4: Illustrations of time series and spatio-temporal data.
In practical applications, spatio-temporal data refers to a collection of observations where each data point is defined by its position in space and time, encapsulating a diverse range of data structures such as graphs, trajectories, and even videos, as noted in [76].
For instance, a spatio-temporal graph (Figure 4 (c)) representing urban traffic flow [286, 287] over time can be understood as spatio-temporal data where each node represents a specific location with specific attributes and edges are weighted by traffic volume, which changes over time. Likewise, trajectory data [281, 288] (Figure 4 (d)) captures the movement of objects through space over time, including their paths, speeds, and changes in direction. Such data is vital for applications in transportation studies, wildlife tracking, and mobile network optimization, where analyzing the patterns of movement and predicting future locations based on historical data are of paramount importance. On this basis, these and other spatio-temporal constructs can be systematically characterized and analyzed.
Building on the definitions provided, we now proceed to succinctly introduce the representative tasks associated with each data category [76].
@@
  • Spatio-Temporal Data Analysis. Spatio-temporal data analysis, while encompassing tasks similar to those in time series analysis, often applies these methodologies within specific application scenarios. For instance, forecasting may focus on traffic flow [43] or air quality [78], utilizing historical data patterns to predict future conditions. Generation tasks might involve creating synthetic trajectories, offering privacy-compliant alternatives to original datasets by replacing sensitive information with generated, anonymized data [81, 290]. Anomaly detection becomes particularly crucial in scenarios such as vehicle trajectory analysis, where deviations from generated normative patterns may indicate unusual or suspicious behaviors [83]. Additionally, spatio-temporal imputation plays a vital role in addressing missing values in multivariate time series [69], ensuring comprehensive and accurate datasets for further analysis. More discussion is in section 5 .
2.2 Why Diffusion Model and Its History
@@
Diffusion models are a class of probability-based generative models. They are called after the mathematical process of diffusion, which is commonly used to describe phenomena such as particle movement in a gas or liquid [1]. In detail, the concept of diffusion models first appeared in statistical physics, used to describe the process of particles moving from areas of high concentration to areas of low concentration [10]. Early diffusion models were primarily concerned with accurately simulating the random diffusion behavior in the generation process.
@@
One of the key breakthroughs in diffusion models occurred in 2015, when researchers proposed a method that combines variational inference to effectively train these models [1]. Since then, the field has experienced rapid development, especially in the area of high-resolution image generation [7]. Since 2020, diffusion models have begun to show their potential in more fields, such as text2image, music generation, and speech synthesis [112, 90, 32]. These advances are due to the optimization of model structures, improvements in training methods, and increased computational resources.
@@
In addition to the field of application, in terms of theory, researchers began to explore how to generate data with specific features by controlling the reverse process. This study direction finally resulted in the development of diffusion models capable of producing high-quality, multidimensional data. In updated computer algorithms, it can be represented as the process of gradually modifying the data distribution, progressively injecting noise until it matches the target distribution, in order to obtain high-quality and realistic synthetic data samples [2].
@@
Recently, more and more researchers and engineers are now focusing their perspectives on the diffusion model, and it has become one of the first choices for generating models [115]. Diffusion models excel at generating high-quality, complex sequences, including time series and spatial-temporal data, with detailed coherence by gradually removing noise. They offer strong control over generations, allowing fine-tuning based on the conditions [114]. These models are flexible and adaptable across various data types and modalities, robust against errors with a gradual noise reduction mechanism, and capable of exploring data diversity for creative outputs. Moreover, they can be integrated with other model types, like autoencoders, to enhance generation quality and control [115].
2.3 Typical Diffusion Models

### v5

@@
Abstract.
Diffusion models have been widely used in time series and spatio-temporal data, enhancing generative, inferential, and downstream capabilities. These models are applied across diverse fields such as healthcare, recommendation, climate, energy, audio, and traffic. By separating applications for time series and spatio-temporal data, we offer a structured perspective on model category, task type, data modality, and practical application domain. This study aims to provide a solid foundation for researchers and practitioners, inspiring future innovations that tackle traditional challenges and foster novel solutions in diffusion model-based data mining tasks and applications. For more detailed information, we have open-sourced a repository [^1] https://github.com/yyysjz1997/Awesome-TimeSeries-SpatioTemporal-Diffusion-Model.
@@ Abstract.
Diffusion Models, Time Series, Spatio-Temporal Data, Generative Model, Temporal Data
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Comput. Surv.; 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Information systems Data mining
1. Introduction
@@
Diffusion models represent a family of probabilistic generative models that undergo optimization through a two-step process involving the injection and subsequent removal of noise across a set of training samples. This process comprises a forward phase, referred to as diffusion, and a reverse phase, known as denoising. By training the model to remove the noise added during the diffusion process, the model learns to generate effective data samples during inference that align closely with the distribution of the training data (Ho et al., 2020; Chang et al., 2023a).
@@
Figure 1. Trends in the cumulative number of papers related to diffusion models for time series and spatio-temporal data.
@@
In recent years, diffusion models have risen to prominence and significantly influenced various domains, including computer vision (CV) (Ho et al., 2020; Dhariwal and Nichol, 2021; Ho et al., 2022; Saharia et al., 2022a, c), natural language processing (NLP) (Austin et al., 2021; Li et al., 2022b; Gong et al., 2022; Lovelace et al., 2024), and general multimodal learning (Ramesh et al., 2022; Saharia et al., 2022b; Yang et al., 2023d; Zhang et al., 2024). This challenges the long-time supremacy of generative adversarial networks (GANs) (Goodfellow et al., 2014; Yang et al., 2023e). Within these areas, diffusion models have demonstrated remarkable capabilities in applications such as text-to-image (Saharia et al., 2022b; Zhang et al., 2023c), instance segmentation (Amit et al., 2021; Wolleb et al., 2022b), 3D shape generation (Vahdat et al., 2022; Nam et al., 2022), molecule design (Xu et al., 2022b; Guo et al., 2023; Abramson et al., 2024), and audio generation (Huang et al., 2022; Kong et al., 2020). Remarkably, diffusion models have also gained popularity as a non-autoregressive alternative for tasks conventionally dominated by autoregressive methods (Chang et al., 2023a). Recently, the introduction of OpenAI Sora (OpenAI Sora, [n. d.]) marked the advent of diffusion models in modeling the physical world embedded within the spacetime continuum, highlighting their critical importance. In addition, AlphaFold 3 (Abramson et al., 2024) proposed by Google DeepMind uses diffusion models to generate 3D atomic coordinates and predict biomolecular structures like proteins, DNA, and RNA.
@@
Temporal data, which primarily includes time series and spatio-temporal data, encapsulates the dynamics of the vast majority of real-world systems (Jin et al., 2023c). These forms of temporal data have been extensively studied and are recognized as crucial for numerous applications (Jin et al., 2023a; Harutyunyan et al., 2019; Jin et al., 2023b). However, deriving universal dynamic laws in the physical world from various data modalities remains a significant challenge within the field. Recently, the area of time series and spatio-temporal modeling has experienced a substantial shift from sensory intelligence towards general intelligence (Jin et al., 2024). This shift is characterized by the emergence of unified foundation models (FMs) that possess versatile temporal data analytical capabilities (Jin et al., 2023c, 2024), challenging the supremacy of domain-specific models. Diffusion models have achieved state-of-the-art results in many modalities, including images, speech, and video (Khanna et al., 2023). Benefiting from the vast and diverse available data in these fields, diffusion models often serve as generative FMs alongside large language models (LLMs) or other foundation models, facilitating rapid development in these areas (Rombach et al., 2022; Ho et al., 2022). In recent years, there has also been an increasing number of diffusion models crafted for modeling time series and spatio-temporal data (Figure 1 ). Also, we have become aware of an increasing number of attempts to use diffusion models for temporal modeling (see Table 1 ). Observing the success of diffusion models, an intriguing question arises: what kind of sparks will emerge from the intersection of time series/spatio-temporal data analysis and diffusion models?
Figure 2. An overview of diffusion models for time series and spatio-temporal data analysis. In the diffusion process, $x_{k}$ and $x_{k-1}$ denote the results after adding noise at step $k$ and k1k-1, respectively. This process can be represented by the size of the controlling steps $\beta_{k}\in(0,1)$, the identity matrix $\mathbf{I}$, and a Gaussian distribution $\mathcal{N}(x;\mu,\sigma)$ of $x$ with the mean $\mu$ and the covariance $\sigma$. During the denoising process, the model attempts to iteratively learn the data distribution by modelling the distribution ${p}_{\theta}(x_{k-1}|x_{k})$. The functions $\mu_{\theta}(\cdot)$ and variance $\sigma_{\theta}(\cdot)$ are the model learnable parameters.
@@
Time series and spatio-temporal data analysis fundamentally rely on a profound understanding of their inherent temporal dynamics, wherein primary tasks predominantly focus on the generative capabilities of backbone models, such as forecasting (Rasul et al., 2021; Li et al., 2022a; Feng et al., 2024), imputation (Tashiro et al., 2021; Wang et al., 2023d; Yun et al., 2023; Wang et al., 2025a) and generation (Narasimhan et al., 2024; Zhang et al., 2024). These analyses center on generating temporal data samples for specific purposes in conditional or unconditional manners. Having witnessed the recent development of time series and spatio-temporal foundation models (Jin et al., 2023c; Liang et al., 2024), whether built upon LLMs or trained from scratch, their success can be attributed to the ability to estimate the distribution of training samples where effective data representations can be drawn. In this regard, diffusion models emerge as a powerful generative framework that enables (1) the modeling of complex patterns within temporal data and (2) the support of a wide range of downstream tasks, as depicted in Figure 2 .
@@
To generate valid data samples for specific tasks, time series and spatio-temporal diffusion models usually operate in an unconditional manner without the need for supervision signals. Given the partially-observed nature of real-world applications (Wang et al., 2024a), conditional diffusion models have emerged. They leverage data labels (e.g., instructions, metadata, or exogenous variables) to regulate the generation process, thereby enabling effective cross-modal prompting that leads to more tailored and improved outcomes (Narasimhan et al., 2024). We present a roadmap in Figure 3 . By training on large-scale temporal data, diffusion models effectively fill the gap in time series/spatio-temporal data generation and exhibit significant potential in solving the puzzle of next-generation, LLM-empowered temporal data-centric agents (Wu et al., 2023; Jin et al., 2024).
@@
Figure 3. Representative diffusion models for time series and spatio-temporal data in recent years.
@@
Despite the promising prospects and rapid advancement of diffusion models in handling time series and spatio-temporal data, there has been a conspicuous lack of systematic analysis of this model family in the existing literature. This article aims to bridge this gap by providing a forward-looking review that elucidates both the ’why’ and the ’how’, detailing the reasons diffusion models are suited for these data modalities and unveiling the mechanisms through which they confer advantages. In this survey, we offer a detailed categorization, engage in thorough reviews, and identify burgeoning trends within this rapidly evolving landscape. Our main contributions are summarized as follows.
@@
2. Background
This paper primarily reviews recent improvements in using diffusion models to solve time series and spatio-temporal data challenges. In this section, we will first define time series and spatio-temporal data, as well as their corresponding tasks in various fields. Then, we will introduce the history of the diffusion model and its advantages. Finally, some different kinds of diffusion models and their variants based on theoretical formula derivations and comparisons with other generative models will be presented.
2.1. Overview of Time Series and Spatio-Temporal Data
Temporal data, particularly time series and spatio-temporal data, are important data structures for a wide range of real-world applications (Jin et al., 2023c). A time series is defined as a sequential arrangement of data points, categorized by their temporal order. These sequences may be univariate, involving a single variable over time, or multivariate, incorporating multiple variables. For instance, daily air quality measurements in a city constitute a univariate time series, while combining daily temperature and humidity readings generates a multivariate series. In our discussions, we employ bold uppercase letters (e.g., $\mathbf{X}$) to represent matrices, bold lowercase (e.g., $\mathbf{x}$) for vectors, calligraphic uppercase (e.g., $\mathcal{X}$) for sets, and standard lowercase (e.g., $x$) for scalars.
We base our formal definitions of time series on those provided in (Jin et al., 2023c, a). Specifically, a univariate time series (Figure 4 (a)), denoted as $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T}$, consists of a sequence of $T$ data points arranged chronologically, where each $x_{t}\in\mathbb{R}$ represents the series’ value at time $t$. Conversely, a multivariate time series (Figure 4 (b)), represented by $\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T})\in\mathbb{R}^{T\times D}$, encompasses a sequence of $T$ data points also in temporal sequence but across $D$ feature channels, with $\mathbf{x}_{t}\in\mathbb{R}^{D}(1\leq t\leq T)$ indicating the series’ values at time $t$ across $D$ different channels. For a comprehensive exploration of time series, we direct the reader to (Jin et al., 2023a).
@@ 2.1. Overview of Time Series and Spatio-Temporal Data
Figure 4. Illustrations of time series and spatio-temporal data.
In practical applications, spatio-temporal data refers to a collection of observations where each data point is defined by its position in space and time, encapsulating a diverse range of data structures such as graphs, trajectories, and even videos, as noted in (Jin et al., 2023c).
For instance, a spatio-temporal graph (Figure 4 (c)) representing urban traffic flow (Sun et al., 2020; Liang et al., 2021a) over time can be understood as spatio-temporal data where each node represents a specific location with specific attributes and edges are weighted by traffic volume, which changes over time. Likewise, trajectory data (Liang et al., 2021b; Chen et al., 2024c) (Figure 4 (d)) captures the movement of objects through space over time, including their paths, speeds, and changes in direction. Such data is vital for applications in transportation studies, wildlife tracking, and mobile network optimization, where analyzing the patterns of movement and predicting future locations based on historical data are of paramount importance. On this basis, these and other spatio-temporal constructs can be systematically characterized and analyzed.
Building on the definitions provided, we now proceed to succinctly introduce the representative tasks associated with each data category (Jin et al., 2023c).
@@
  • Spatio-Temporal Data Analysis. Spatio-temporal data analysis, while encompassing tasks similar to those in time series analysis, often applies these methodologies within specific application scenarios. For instance, forecasting may focus on traffic flow (Wen et al., 2023) or air quality (Chen et al., 2024b), utilizing historical data patterns to predict future conditions. Generation tasks might involve creating synthetic trajectories, offering privacy-compliant alternatives to original datasets by replacing sensitive information with generated, anonymized data (Zhu et al., 2024a, b). Anomaly detection becomes particularly crucial in scenarios such as vehicle trajectory analysis, where deviations from generated normative patterns may indicate unusual or suspicious behaviors (Li et al., 2024a). Additionally, spatio-temporal imputation plays a vital role in addressing missing values in multivariate time series (Yun et al., 2023; Du et al., 2024), ensuring comprehensive and accurate datasets for further analysis. More discussion is given in section 5 .
2.2. Why Diffusion Model and Its History
@@
Diffusion models are a class of probability-based generative models. They are called after the mathematical process of diffusion, which is commonly used to describe phenomena such as particle movement in a gas or liquid (Sohl-Dickstein et al., 2015). In detail, the concept of diffusion models first appeared in statistical physics, used to describe the process of particles moving from areas of high concentration to areas of low concentration (Jarzynski, 1997). Early diffusion models were primarily concerned with accurately simulating the random diffusion behavior in the generation process.
@@
One of the key breakthroughs in diffusion models occurred in 2015, when researchers proposed a method that combines variational inference to effectively train these models (Sohl-Dickstein et al., 2015). Since then, the field has experienced rapid development, especially in the area of high-resolution image generation (Rombach et al., 2022). Since 2020, diffusion models have begun to show their potential in more fields, such as text2image, music generation, and speech synthesis (Saharia et al., 2022b; Wang, 2023; Richter et al., 2023). These advances are due to the optimization of model structures, improvements in training methods, and increased computational resources.
@@
In addition to the field of application, in terms of theory, researchers began to explore how to generate data with specific features by controlling the reverse process. This study direction finally resulted in the development of diffusion models capable of producing high-quality, multidimensional data. In updated computer algorithms, it can be represented as the process of gradually modifying the data distribution, progressively injecting noise until it matches the target distribution, in order to obtain high-quality and realistic synthetic data samples (Ho et al., 2020).
@@
Recently, more and more researchers and engineers are now focusing their perspectives on the diffusion model, and it has become one of the first choices for generating models (Yang et al., 2023e). Diffusion models excel at generating high-quality, complex sequences, including time series and spatio-temporal data, with detailed coherence by gradually removing noise. They offer strong control over generations, allowing fine-tuning based on the conditions (Croitoru et al., 2023). These models are flexible and adaptable across various data types and modalities, robust against errors with a gradual noise reduction mechanism, and capable of exploring data diversity for creative outputs. Moreover, they can be integrated with other model types, like autoencoders, to enhance generation quality and control (Yang et al., 2023e).
2.3. Typical Diffusion Models